# A Hitchhikers Guide to Fine-Grained Face Forgery Detection Using Common Sense Reasoning

### v1

@@ Abstract
Explainability in artificial intelligence is crucial for restoring trust, particularly in areas like face forgery detection, where viewers often struggle to distinguish between real and fabricated content. Vision and Large Language Models (VLLM) bridge computer vision and natural language, offering numerous applications driven by strong common-sense reasoning. Despite their success in various tasks, the potential of vision and language remains underexplored in face forgery detection, where they hold promise for enhancing explainability by leveraging the intrinsic reasoning capabilities of language to analyse fine-grained manipulation areas. As such, there is a need for a methodology that converts face forgery detection to a Visual Question Answering (VQA) task to systematically and fairly evaluate these capabilities. Previous efforts for unified benchmarks in deepfake detection have focused on the simpler binary task, overlooking evaluation protocols for fine-grained detection and text-generative models. We propose a multi-staged approach that diverges from the traditional binary decision paradigm to address this gap. In the first stage, we assess the models’ performance on the binary task and their sensitivity to given instructions using several prompts. In the second stage, we delve deeper into fine-grained detection by identifying areas of manipulation in a multiple-choice VQA setting. In the third stage, we convert the fine-grained detection to an open-ended question and compare several matching strategies for the multi-label classification task. Finally, we qualitatively evaluate the fine-grained responses of the VLLMs included in the benchmark. We apply our benchmark to several popular models, providing a detailed comparison of binary, multiple-choice, and open-ended VQA evaluation across seven datasets. https://nickyfot.github.io/hitchhickersguide.github.io/
@@ 1 Introduction
Numerous successful works on deepfake detection have been proposed in the literature to tackle the risks of face forgery [37, 76, 52, 7, 80]. Existing methods primarily rely on deep binary classifiers, resulting in black-box models that predict whether an input is real or fake. Consequently, explaining why those predictions are being made is not straightforward. To handle this issue, a few interpretable deepfake detection methods have been introduced by examining attention maps or weight activations [86, 71] to identify fine-grained areas of manipulation; however, these are based on a post-hoc analysis and thereby do not intrinsically incorporate an explainable mechanism. On the other hand, Vision Large Language Models (VLLMs) have emerged as a pioneering branch of generative Artificial Intelligence (AI), showcasing advancements in common sense reasoning and an inherent explainability that arises from the intrinsic nature of language [22]. They have demonstrated impressive capabilities in tasks such as Visual Question Answering (VQA) [47] and the generation of descriptive content for downstream applications [70], hence bridging the gap between vision-language understanding and contextual reasoning. However, despite these achievements, the explainable power of VLLMs remains under-explored in the field of deepfake detection, with only a handful of works exploring the vision and language capabilities for the binary classification of fake/real images [26, 8, 31, 69] all of which are evaluated on different benchmarks and metrics.
@@
Explainable fine-grained detection –that is, identifying manipulation beyond the binary fake/real decision– in natural language is still in its infancy. However, as VLLM works for deepfake detection are expected to appear following the overwhelming success of foundation models in other tasks, two research questions need to be addressed: 1) “To what extent can existing VLLMs detect deepfake images and what rationale supports the decision?” and 2) “How do we fairly and comprehensively evaluate VLLMs in the fine-grained task?”. In deepfake detection, benchmarks have mainly focused on binary decisions and discriminative networks [81], making them unsuitable to answer these research questions. Indeed, they do not propose a unified method to match the generated responses to fine-grained categories. Similarly, existing benchmarks in Visual Question Answering (VQA) [19, 10] primarily address multi-class tasks, which may not be suitable for the multi-label nature of fine-grained deepfake detection as highlighted in [63].
@@
In this work, our objective is to conduct a thorough quantitative and qualitative evaluation of VLLMs for the task of fine-grained deepfake detection in a systematic and scientific approach, employing a multi-stage protocol. In the first stage, we assess the models’ performance on the binary task using various prompts while also evaluating the model’s sensitivity to the provided instructions. In the second stage, we delve into fine-grained detection, aiming to identify areas of manipulation within a closed-form Visual Question Answering (VQA) framework, i.e. what areas from a given list are identified as manipulated. Subsequently, in the third stage, we extend our investigation by converting the fine-grained detection task into an open-ended question –that is, identifying areas without instructing the model to select from a list of categories. Here, as the task is a multi-label problem, we compare two matching strategies: a) using the cosine similarity between the generated text and ground truth labels and b) counting the occurrence of the class name in the generated text. Finally, we qualitatively evaluate the fine-grained responses generated by the VLLMs included in our benchmark, providing nuanced and new insights into their performance.
@@
  • We introduce a novel approach in deepfake detection by proposing a method to transform the discriminative task into a Visual Question Answering (VQA) multi-label problem. To the best of our knowledge, this is the first work to do so in the field of deepfake detection, offering a fresh perspective on explainability through fine-grained analysis.
  • We establish a systematic, unified evaluation framework for current state-of-the-art (SOTA) VLLMs, facilitating consistent assessment across different models. This framework is designed to be extendable, ensuring fair and comprehensive comparisons with future models, thus promoting transparency and reproducibility in the evaluation process.
@@ 2 Related Work
Benchmarks for classification tasks [19, 38, 2] in a VQA setting typically address the multi-class paradigm, which may not be appropriate for addressing explainability in DeepFake detection by adopting a multi-label fine-grained strategy, as several areas can be manipulated at once. A few preliminary works in DeepFake detection [3, 26] benchmark ChatGPT4111https://openai.com/gpt-4 and Gemini [^2] https://gemini.google.com; however, these have primarily focused on the more straightforward binary task and did not explore the reasoning capabilities of VLLMs for fine-grained labels. Furthermore, both these works focused on VLLMs that are not open-sourced with limited information regarding their training set and architecture; thus, it is not possible to assess whether the benchmarks are, in fact, zero-shot or whether they have been trained on deepfake-related image-language pairs. Within DeepFake detection, the vast majority of benchmarking works [60, 65, 46, 81, 34] have focused on binary discriminative networks and are therefore not fit to evaluate the capabilities of generative models such as VLLM, particularly for fine-grained labels.
@@ 4.1 Tested VLLMs
We select four open-source state-of-the-art VLLMs to be included in this benchmark; specifically, we include LlaVa-1.5 [43] (an improved version of the LlaVa architecture [44]), BLIP2 [35] and finally InstructBlip [15] with Flan-T5 and Flan-T5-xxl language generators [13]. Finally, for the binary task, we include the CLIP [59] performance as a baseline and compare it against an ensemble of BLIP2 [35] and LlaVa-1.5 [43] following the ensembling strategies for VQA tasks [5], and GPT4v as an upper bound [^4] https://openai.com/index/gpt-4v-system-card/. Experiments using GPT4v are performed on a subset of 5k samples selected from each dataset and thus the results may be susceptible to some degree of bias from the sampling, which needs to be considered when comparing the models. The selection of the VLLM is guided by three factors. First, we select architectures with publicly available weights and training methods to ensure transparency and fairness in the evaluation. Second, we include methods that generate output in Natural Language rather than a set of features or classification predictions. Finally, we select methods that have achieved state-of-the-art performance on several zero-shot tasks. Additional model details, such as the number of parameters and pre-training information of the tested models, can be found in Appendix A.
@@ 5.1 Binary Classification
Overall model performance: We average the performance of the three best-performing synonyms on all nine benchmarks in  Fig. 3. No model clearly outperforms others across all metrics and datasets; however, we can observe competitive performance from BLIP-2 [35] on the binary task, even though it is the smallest model in terms of parameters. We also see that all models struggle with the more challenging in-the-wild datasets, such as CelebDF [41], which highlights the need for further development to achieve adequate generalisation.
@@ 5.3 Qualitative Evaluation
Overall performance: In the simpler binary setting, BLIP-2 is more robust to instruction than other models with more parameters; however, when it comes to fine-grained evaluation, larger models show an advantage in reasoning and identifying areas of manipulation in the open-ended and multiple-choice settings. It is, however, worth noting that no model clearly outperforms others across all metrics and datasets. All of the results presented are based on zero-shot evaluations, where models are tested without being specifically trained for deepfake detection. Despite this, the models are able to leverage a semantic mapping between language and visual input from their very vast pre-training, giving them an inherent concept of "real" versus "fake." This capability suggests that these models possess some degree of understanding when it comes to identifying deepfakes. However, this general understanding falls far behind that of task-specific models. When we fine-tune the vision encoder, there is a notable improvement in performance. The vision-language models can better capture details and nuances in the input data, which enhances their deepfake detection capabilities. Nevertheless, due to the scarcity of high-quality captions and large-scale vision-language datasets tailored to deepfake detection, the improvements remain limited and only in the binary task. Overall, addressing these limitations—by creating specialised datasets and foundation models—could lead to substantial advancements in this area.
Limitations and Future Work: As the models in this work are all evaluated under zero-shot settings, their performance is below that of purpose-build networks seen in previous works [81], particularly for more challenging in-the-wild datasets. This further highlights the need for task specific models and more fine-grained deepfake datasets, which is a key finding of the experiments conducted in this work. A significant limitation is the lack of detailed language descriptions in datasets, making qualitative evaluation harder. Additionally, current datasets lack fine-grained labels, restricting assessments of manipulations to pseudo-fakes and SeqDeepFake [63]. Furthermore, as both the pertaining and evaluation datasets are not unbiased, the performance of all VLLMs is susceptible to the bias of the datasets, which is not addressed in this or previous benchmarks [81]. Identifying these shortcomings is important for future works on the task, particularly as VLLMs gain traction.

### v2

@@ Abstract
Explainability in artificial intelligence is crucial for restoring trust, particularly in areas like face forgery detection, where viewers often struggle to distinguish between real and fabricated content. Vision and Large Language Models (VLLM) bridge computer vision and natural language, offering numerous applications driven by strong common-sense reasoning. Despite their success in various tasks, the potential of vision and language remains underexplored in face forgery detection, where they hold promise for enhancing explainability by leveraging the intrinsic reasoning capabilities of language to analyse fine-grained manipulation areas. For that reason, few works have recently started to frame the problem of deepfake detection as a Visual Question Answering (VQA) task, nevertheless omitting the realistic and informative open-ended multi-label setting. With the rapid advances in the field of VLLM, an exponential rise of investigations in that direction is expected. As such, there is a need for a clear experimental methodology that converts face forgery detection to a Visual Question Answering (VQA) task to systematically and fairly evaluate different VLLM architectures. Previous evaluation studies in deepfake detection have mostly focused on the simpler binary task, overlooking evaluation protocols for multi-label fine-grained detection and text-generative models. We propose a multi-staged approach that diverges from the traditional binary evaluation protocol and conducts a comprehensive evaluation study to compare the capabilities of several VLLMs in this context. In the first stage, we assess the models’ performance on the binary task and their sensitivity to given instructions using several prompts. In the second stage, we delve deeper into fine-grained detection by identifying areas of manipulation in a multiple-choice VQA setting. In the third stage, we convert the fine-grained detection to an open-ended question and compare several matching strategies for the multi-label classification task. Finally, we qualitatively evaluate the fine-grained responses of the VLLMs included in the benchmark. We apply our benchmark to several popular models, providing a detailed comparison of binary, multiple-choice, and open-ended VQA evaluation across seven datasets. https://nickyfot.github.io/hitchhickersguide.github.io/
@@ 1 Introduction
Numerous successful works on deepfake detection have been proposed in the literature to tackle the risks of face forgery [37, 77, 52, 7, 81]. Existing methods primarily rely on deep binary classifiers, resulting in black-box models that predict whether an input is real or fake. Consequently, explaining why those predictions are being made is not straightforward. To handle this issue, a few interpretable deepfake detection methods have been introduced by examining attention maps or weight activations [88, 72] to identify fine-grained areas of manipulation; however, these are based on a post-hoc analysis and thereby do not intrinsically incorporate an explainable mechanism. On the other hand, Vision Large Language Models (VLLMs) have emerged as a pioneering branch of generative Artificial Intelligence (AI), showcasing advancements in common sense reasoning and an inherent explainability that arises from the intrinsic nature of language [22]. They have demonstrated impressive capabilities in tasks such as Visual Question Answering (VQA) [47] and the generation of descriptive content for downstream applications [71], hence bridging the gap between vision-language understanding and contextual reasoning. However, despite these achievements, the explainable power of VLLMs remains under-explored in the field of deepfake detection, with only a handful of works mostly exploring the vision and language capabilities for the binary classification of fake/real images [26, 8, 31, 70] all of which are evaluated on different benchmarks and metrics. To the best of our knowledge, Zhang et al.  [87] is the only work employing a VQA approach in deepfake detection by proposing to extend the FF++ dataset with captions in natural language generated by humans. However, this work targets only one manipulated region at once, while deepfakes can incorporate several stacked manipulations [63, 68]. Moreover, the provided augmented FF++ dataset does not allow for cross-dataset evaluation in a VQA setting without an extensive annotation effort, making it difficult to investigate the generalisation aspect. In addition, the fine-grained evaluation in previous works is limited as the more challenging open-ended VQA task is not explored.
@@
Explainable fine-grained detection –that is, identifying manipulation beyond the binary fake/real decision– in natural language is still in its infancy. However, as VLLM works for deepfake detection are expected to appear following the overwhelming success of foundation models in other tasks, two research questions need to be addressed: 1) “To what extent can existing VLLMs detect deepfake images and what rationale supports the decision?” and 2) “How do we fairly and comprehensively evaluate VLLMs in the fine-grained task?”. In deepfake detection, benchmarks have mainly focused on binary or multi-class decisions and discriminative networks [82, 87], making them unsuitable to answer these research questions. Indeed, they do not propose a unified method to match the generated responses to fine-grained multi-label categories. Similarly, existing benchmarks in Visual Question Answering (VQA) [19, 10] primarily address multi-class tasks, which may not be suitable for the multi-label nature of fine-grained deepfake detection as highlighted in [63].
@@
In this work, our objective is to conduct a thorough quantitative and qualitative evaluation of VLLMs for the task of fine-grained multi-label deepfake detection in a systematic and scientific approach, employing a multi-stage protocol without costly human captioning efforts. In the first stage, we assess the models’ performance on the binary task using various prompts while also evaluating the model’s sensitivity to the provided instructions. In the second stage, we delve into multi-label fine-grained detection, aiming to identify areas of manipulation within a multiple-choice Visual Question Answering (VQA) framework, i.e. what areas from a given list are identified as manipulated. Subsequently, in the third stage, we extend our investigation by converting the fine-grained detection task into an open-ended question –that is, identifying areas without instructing the model to select from a list of categories. Here, as the task is a multi-label problem, we compare two matching strategies: a) using the cosine similarity between the generated text and ground truth labels and b) counting the occurrence of the class name in the generated text. Finally, we qualitatively evaluate the fine-grained responses generated by the VLLMs included in our benchmark, providing nuanced and new insights into their performance.
@@
  • We introduce a novel evaluation protocol for deepfake detection under the Visual Question Answering (VQA) multi-label setting and without the use of human annotations. This is different from  [87] which is based on a succession of yes/no questions for fine-grained areas, resulting in a binary classification setting. It also relies on a relatively small dataset that cannot be extended without costly annotation efforts. In addition, our multi-stage protocol allows for open-ended VQA evaluation, which is a more challenging task. To the best of our knowledge, this is the first work to do so in the field of deepfake detection, offering a fresh perspective on explainability through fine-grained multi-label analysis.
  • We present a systematic, unified evaluation study of current state-of-the-art (SOTA) VLLMs, facilitating consistent assessment across different models. This framework is designed to be extendable to any existing or future deepfake dataset, ensuring fair and comprehensive comparisons with future models, thus promoting transparency and reproducibility in the evaluation process.
@@ 2 Related Work
Benchmarks for classification tasks [19, 38, 2] in a VQA setting typically address the multi-class paradigm, which may not be appropriate for addressing explainability in DeepFake detection by adopting a multi-label fine-grained strategy, as several areas can be manipulated at once. A few preliminary works in DeepFake detection [3, 26, 87] benchmark ChatGPT4111https://openai.com/gpt-4 and Gemini [^2] https://gemini.google.com; however, these have primarily focused on the more straightforward binary task and did not explore the reasoning capabilities of VLLMs for fine-grained labels. Furthermore, both these works focused on VLLMs that are not open-sourced with limited information regarding their training set and architecture; thus, it is not possible to assess whether the benchmarks are, in fact, zero-shot or whether they have been trained on deepfake-related image-language pairs. Zhang et al.  [87] propose extending FF++ annotations with captions in natural language using human annotators. However, this method is limited to binary decisions, while a given deepfake image can undergo several manipulations. Furthermore, it does not explore the open-ended VQA setting and does not offer a method for cross-dataset evaluation without the need for a costly annotation process. Within DeepFake detection, the vast majority of benchmarking works [60, 65, 46, 82, 34] have focused on binary discriminative networks and are therefore not fit to evaluate the capabilities of generative models such as VLLM, particularly for fine-grained labels.
@@ 4.1 Tested VLLMs
We select four open-source state-of-the-art VLLMs to be included in this benchmark; specifically, we include LlaVa-1.5 [43] (an improved version of the LlaVa architecture [44]), BLIP2 [35] and finally InstructBlip [15] with Flan-T5 and Flan-T5-xxl language generators [13]. Finally, for the binary task, we include the CLIP [59] performance as a baseline and compare it against an ensemble of BLIP2 [35] and LlaVa-1.5 [43] following the ensembling strategies for VQA tasks [5], and GPT4v as an upper bound [^4] https://openai.com/index/gpt-4v-system-card/. Experiments using GPT4v are performed on a subset of 5k samples selected from each dataset, and thus, the results may be susceptible to some degree of bias from the sampling, which needs to be considered when comparing the models. The selection of the VLLM is guided by three factors. First, we select architectures with publicly available weights and training methods to ensure transparency and fairness in the evaluation. Second, we include methods that generate output in Natural Language rather than a set of features or classification predictions. Finally, we select methods that have achieved state-of-the-art performance on several zero-shot tasks. Additional model details, such as the number of parameters and pre-training information of the tested models, can be found in Appendix A.
@@ 5.1 Binary Classification
Overall model performance: We average the performance of the three best-performing synonyms on all nine benchmarks in  Fig. 3. No model clearly outperforms others across all metrics and datasets; however, we can observe competitive performance from BLIP-2 [35] on the binary task, even though it is the smallest model in terms of parameters. We also see that all models struggle with the more challenging in-the-wild datasets, such as CelebDF [41], which highlights the need for further development to achieve adequate generalisation. Performance of GPT4v should be treated as an upper bound as we cannot assess whether the model has been trained on samples of the selected datasets. We see that GPT4v vastly outperforms the selected VLLMs on three benchmarks and has comparable performance on the rest, with the exception of FF++.
@@ 5.3 Qualitative Evaluation
Overall performance: In the simpler binary setting, BLIP-2 is more robust to instruction than other models with more parameters; however, when it comes to fine-grained evaluation, larger models show an advantage in reasoning and identifying areas of manipulation in the open-ended and multiple-choice settings. It is, however, worth noting that no model clearly outperforms others across all metrics and datasets. All of the results presented are based on zero-shot evaluations, where models are tested without being specifically trained for deepfake detection. Despite this, the models are able to leverage a semantic mapping between language and visual input from their very vast pre-training, giving them an inherent concept of “real” versus “fake”. This capability suggests that these models possess some degree of understanding when it comes to identifying deepfakes. However, this general understanding falls far behind that of task-specific models. When we fine-tune the vision encoder, there is a notable improvement in performance. Vision-language models can better capture details and nuances in input data, which enhances their deepfake detection capabilities. Nevertheless, due to the scarcity of high-quality captions and large-scale vision-language datasets tailored to deepfake detection, the improvements remain limited and only in the binary task. Overall, addressing these limitations—by creating specialised datasets and foundation models—could lead to substantial advancements in this area.
Limitations and Future Work: As the models in this work are all evaluated under zero-shot settings, their performance is below that of purpose-build networks seen in previous works [82], particularly for more challenging in-the-wild datasets. This further highlights the need for task-specific models and more fine-grained deepfake datasets, which is a key finding of the experiments conducted in this work. A significant limitation is the lack of detailed language descriptions in datasets, making qualitative evaluation harder. Additionally, current datasets lack fine-grained labels, restricting assessments of manipulations to pseudo-fakes and SeqDeepFake [63]. Furthermore, as both the pertaining and evaluation datasets are not unbiased, the performance of all VLLMs is susceptible to the bias of the datasets, which is not addressed in this or previous benchmarks [82]. Identifying these shortcomings is important for future works on the task, particularly as VLLMs gain traction.